# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v2

@@ Abstract.
Answering natural language (NL) questions about tables, known as Tabular Question Answering (TQA), is crucial because it allows users to quickly and efficiently extract meaningful insights from structured data, effectively bridging the gap between human language and machine-readable formats. Many of these tables are derived from web sources or real-world scenarios, which require meticulous data preparation (or data prep) to ensure accurate responses. However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation. This question-aware data preparation involves specific tasks such as column augmentation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Executes the generated code to process the table. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the state-of-the-art TQA solutions through question-aware data preparation.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024b) rely on large language models (LLMs), either utilizing LLMs as black-boxes or leveraging them for code generation.
(a) WikiTQ
(b) TabFact
@@
Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets, with statistics calculated from a sample of 500 instances from each dataset.
(a) Question-Aware Column Augmentation: Necessary information for QA is not explicitly available in tables.
(b) Question-Aware Value Normalization: Heterogeneous data representations or formats in tables, and how to normalize is query-dependent.
(c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
Figure 2. Examples of Question-Aware Data Preparation.
@@
Question-Aware Data Preparation. As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers. Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The results show that 84% and 76% of the errors arise from insufficiently addressing the data problems at various granularities (table-level, column-level, and cell-level) in relation to the NL questions. This highlights the critical role of thorough data prep in ensuring accurate responses for TQA.
Figure 2 shows several typical data problems critical to the quality of TQA, which in turn motivate the need of question-aware data preparation.
Example 0.
(a) Column-Level: Missing Column. The example in Figure 2(a) shows the Country column that the NL question $Q_{1}$ requires is not explicitly available.
In this case, an appropriate data prep could address this problem by extracting country information from Cyclist, i.e., augmenting the tables with a new Country column. Clearly, this data augmentation is driven by the specific need of the individual query.
(b) Cell-Level: Inconsistent Data Types or Formats. Figure 2(b) highlights the inconsistency among the data in different columns (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$).
This essentially demands data normalization. This normalization has to be question-aware, because different questions may require different types of normalization. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Table-Level: Irrelevant Columns. As shown in Figure 2(c), some columns in the tables are irrelevant to the questions. For example, $Q_{7}$ does not need the record column in table $T_{7}$. On the other hand, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
Filtering out the columns irrelevant to a question is important to TQA, as these irrelevant columns add complexity, potentially misleading LLMs to generate wrong answers. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may confuse the LLM when answering $Q_{7}$. Moreover, this column filtering, avoiding feeding all columns to the LLM, effectively reduces LLM inference cost.
The above examples demonstrate the urgent need of data prep to ensure the quality of TQA. However, blindly applying the existing data prep techniques to address all possible data problems tends to be neither effective nor efficient due to overlooking the specific needs of different NL questions.
In this work, we thus propose to study a new problem, namely question-aware data preparation. Unlike the conventional data prep which prepares tables once at an offline stage to serve many incoming analytics queries (questions), question-aware data prep addresses question specific data problems at online query stage. Once receiving a question, it tailors the tables to the specific need of this question, e.g., performing tasks such as question-specific column augmentation and filtering, as well as question-aware value normalization or conversion.
@@
Key Challenges. Building a system to support question-aware data prep is challenging because of the diversity of both the data problems the tables encounter and the questions. As illustrated in the above examples, different questions may demand different types of data prep, e.g., augmentation, normalization, or filtering. Furthermore, even if facing the same type of data prep problem, different questions may require different ways to handle it. For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers. Existing LLM-based TQA solutions do not explicitly consider the question-aware data prep problem, thus suffering from the data issues mentioned above. For a detailed analysis, please see Section 2.
@@
AutoPrep: A Hierarchical, Multi-Stage Approach. To address these challenges, we propose AutoPrep, which features two key ideas. First, inspired by the design principles of modern DBMS, in particular the distinction between the logical operations and their physical implementations, AutoPrep separates the high level, logical data prep operations and the concrete data prep methods. That is, it introduces a planning stage to generate a logical plan for each question. The logical plan is composed of a sequence of high-level data prep operations needed by the corresponding questions, including column augmentation, value normalization, or filtering irrelevant data, as shown in Figure 2. AutoPrep then introduces another stage to map the selected logical operations to the corresponding physical operations. In this way, AutoPrep decomposes a data prep task into smaller sub-tasks which are easier to solve than the original complex task. Moreover, it makes AutoPrep extensible, allowing us to easily develop new specialized implementations w.r.t. each type of data prep operations, or supporting new operation types.
@@
Second, unlike conventional DBMS, in AutoPrep the physical operators w.r.t. each logical operator are not developed beforehand. Instead, given a data prep operation in the logical plan, AutoPrep generates a physical implementation specialized to the question on the fly. Taking the question specific need into consideration, this customized implementation thus tends to meet the unique need of different questions.
@@
Driven by the above insight, we design the system into a multi-stage architecture.
@@
The Planning Stage: Unlike DBMS where the logical operators are already available beforehand via SQL queries, AutoPrep has to decide on the appropriate logical data prep operations and in turn form the logical plan by analyzing the NL question and the given table. This is done during the planning stage.
@@
The Programming Stage: This stage translates the logical plan into a physical plan by generating low-level executable code, which is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats, etc.
@@
The Executing Stage: This stage executes the generated code for each operation and returning any errors encountered to the Programming stage for debugging.
@@
We implements this AutoPrep architecture with the popular LLM-based Multi-Agent framework, which uses multiple small, independent agents to collaboratively solve complex problems.
@@
More specifically, we design a Planner agent. Corresponding to the Planning Stage, it suggests a tailored sequence of high-level operations that meet the specific needs of these questions, leveraging the semantics understanding and reasoning ability of LLMs. The key technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer and guiding the agent’s reasoning based on this sketch. Compared with the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the relationships between questions and tables.
@@
AutoPrep also contains a set of Programmer agent, each of which synthesizes a question specific implementation w.r.t. one logical data prep operation. However, existing LLM-based code synthesis often produces overly generic code that fails to effectively address the heterogeneity challenges of tables. For example, values may have diverse syntactic formats (e.g., “19-Oct” and “9/14” in $T_{6}$) or or semantic representations (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code tailored to such variations. To address this, we develop a tool-augmented approach that improves the LLM’s code generation capabilities by utilizing predefined API functions. This approach effectively resolves the issue of overly generic code by allowing the LLM to leverage specialized functions tailored to variations in table values. Moreover, corresponding to the Executing stage, we design an Executor agent that executes the generated code to process the table.
@@ 2.1. Tabular Question Answering
Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) ${A_{1},A_{2},\ldots,A_{m}}$ and $n$ rows ${r_{1},r_{2},\ldots,r_{n}}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table. The problem of tabular question answering (TQA) is to generate an answer, denoted as $A$, in response to question $Q$ based on the information in table $T$. By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
@@
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
@@ 2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to specific needs of question $Q$.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
  • Augment: a data prep task that augments a new column for table $T$ based on existing columns, to better answer $Q$. This task typically involves column combinations via arithmetic operations, value extraction and inference, etc.
  • Normalize: a data prep task that normalize types or formats of the values in a column of $T$ based on the needs of $Q$. This task typically involves value representation or format normalization, type conversion, etc.
  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answering question $Q$. This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
@@
(a) The Augment operation: Figure 2(a) shows three examples of Augment operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
(b) The Normalization operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing country formats in $T_{4}$, converting strings into integers in $T_{5}$, and normalizing value formats in Date in $T_{6}$. We can see that, to implement a specific operation (e.g., in $T_{4}$), we sometimes need to call external tools, like a function clean_country from an external library.
(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{5}$ and $T_{6}$. The Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, potentially misleading the LLM. Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
The above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, due to the separation of logical operations and physical implementations in AutoPrep, the set of data prep operations can be easily extended to support new data prep tasks.
@@
Example 0.
To better answer the example question $Q_{1}$ over $T_{1}$ in Figure 2(a), we first generate an Augment operation to transform $T_{1}$ into $T^{\prime}_{1}$ with three columns: Cyclist, Medal, and Country. Next, we generate a Normalize operation, similar to the one in $T_{4}$, to standardize the values in Country into a unified format, producing table $T^{\prime\prime}_{1}$. Finally, we leverage LLMs, either as black-box or for code generation, to compute the answer $A_{1}$ to $Q_{1}$ from the prepared table $T^{\prime\prime}_{1}$.
2.3. Limitations of Existing LLM-Based Solutions

### v3

@@ Abstract.
Answering natural language (NL) questions about tables, known as Tabular Question Answering (TQA), is crucial because it allows users to quickly and efficiently extract meaningful insights from structured data, effectively bridging the gap between human language and machine-readable formats. Many of these tables are derived from web sources or real-world scenarios, which require meticulous data preparation (or data prep) to ensure accurate responses. However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation. This question-aware data preparation involves specific tasks such as column derivation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Executes the generated code to process the table. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the state-of-the-art TQA solutions through question-aware data preparation.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) rely on large language models (LLMs).
@@
Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets.
@@
As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers. Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019) (More details of the error analysis can be found in our technical report (tec, 2025)). The results indicate that 84% and 76% of the errors stem from inadequately addressing data prep issues, including missing semantics, inconsistent values, and irrelevant columns, as illustrated as follows.
(a) Missing Semantics: Required semantic information for NL questions is not explicitly present in the tables.
(b) Value Inconsistency: Heterogeneous data representations or formats across table values, where normalization is question-dependent.
(c) Irrelevant Columns: Presence of unnecessary columns that increase complexity and induce errors for LLMs.
Figure 2. Examples of Data Preparation Issues for Table Question Answering.
(1) Missing Semantics. This data prep issue arises when a table lacks the necessary semantics to address the specific requirements of the NL question. That is, although some columns in the table may be related to the question, they do not directly provide the required semantic information. As shown in Figure 2(a), the semantics needed for the NL questions, such as country and GDP, are not explicitly present in the tables. Therefore, to ensure accurate responses in TQA, it is essential to perform column derivation from existing columns, such as extracting country information from text or applying mathematical operations across multiple columns.
(2) Inconsistent Values. This data preparation issue arises when the values within a particular column are inconsistent across different records due to varying data representations (e.g., inconsistent date formats), as illustrated in Figure 2(b). This situation necessitates column normalization, which must be question-aware, as different questions may require different types of normalization. For instance, as $Q_{5}$ requires comparing numbers in $T_{5}$ but the corresponding columns are of string type, it is necessary to convert the strings in fraction format into numerical values.
(3) Irrelevant Columns. In some tables, although many columns are available, only a few are actually relevant to the NL question. For example, as shown in Figure 2(c), $Q_{7}$ does not require the Summary column in table $T_{7}$. Similarly, in $T_{8}$, although there are 12 columns, only 2 are needed to answer $Q_{8}$. Thus, column filtering, i.e., retaining only the columns directly related to the question, is crucial for TQA, as irrelevant columns add complexity and can mislead LLMs into producing incorrect answers. For instance, the Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column.
This highlights the critical role of thorough data prep in ensuring accurate TQA. Note that, although issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks. Nevertheless, we recognize their significance and leave the exploration of broader data prep challenges to future work.
Question-Aware Data Preparation. In this paper, we thus propose to study a new problem, namely question-aware data preparation. The novelty of this problem lies in the need to deeply understand the semantics of the NL question in order to guide data prep operations over the structured table. Unlike traditional data prep, which is performed offline and independent of downstream tasks, question-aware data prep is conducted online and must dynamically tailor the table to the specific demands of the question. This introduces new challenges due to the diversity and ambiguity of NL questions and the heterogeneity of tabular data, requiring precise semantic alignment between the question and the structured table. For example, as shown in Figure 2(a), given question $Q_{1}$ that requests country-specific information from a text column, table $T_{1}$ needs to be transformed to extract and create a new Country column, which may not be considered in traditional data prep.
@@
However, building a system to support question-aware data prep is challenging due to the semantic alignment between the table and the NL question. Specifically, different questions may demand different types of data prep, e.g., derivation, normalization, or filtering. Furthermore, even if facing the same type of data prep issue, different questions may require different ways to handle it. For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers. Although LLMs show promise for interpreting semantics in NL questions, simply applying them to address all potential data prep issues has shown ineffective, often leading to false negatives and false positives, as illustrated in Section 2.3.
@@
AutoPrep: A Hierarchical, Multi-Stage Approach. To address the challenges, we propose AutoPrep, which features two key ideas. First, drawing inspiration from modern DBMS, particularly the distinction between logical operations and their physical implementations, AutoPrep separates high-level, logical data prep operations from the concrete methods used for execution. Specifically, it introduces a planning stage that generates a logical plan for each question, consisting of a sequence of high-level data prep operations tailored to the question’s needs, such as column derivation, normalization and filtering shown in Figure 2. In the next stage, AutoPrep maps these logical operations to the corresponding physical implementations. This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process more easier to solve. Moreover, this modular design makes AutoPrep extensible, enabling the development of specialized implementations for each type of data prep operation or the introduction of new operation types.
@@
Second, unlike conventional DBMSs, AutoPrep does not predefine physical operations for each logical operation. Instead, for each data prep operation in the logical plan, AutoPrep generates a physical implementation that is specialized to the specific question on the fly. By considering the unique requirements of the NL question, this customized implementation ensures that the data preparation is closely aligned with the needs of different questions.
@@
Building on the above insights, we design the AutoPrep system with a multi-stage architecture.
@@
  • The Planning Stage: Unlike traditional DBMS, where logical operations are already available beforehand via SQL queries, AutoPrep has to determine the appropriate logical data prep operations by analyzing the semantic alignment between the table and the NL question. This process occurs during the planning stage, where the logical plan is formed.
@@
  • The Programming Stage: This stage converts the logical plan into a physical plan by generating low-level executable code, selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, and customizing the code to align the table’s structure with the NL question’s semantics.
@@
  • The Executing Stage: This stage executes the generated code for each operation and returns any errors encountered to the Programming stage for debugging.
@@
We implement AutoPrep using the popular LLM-based Multi-Agent framework  (Han et al., 2024), which leverages multiple small, independent agents working collaboratively to solve complex problems.
@@
More specifically, we design a Planner agent, which corresponds to the Planning Stage and suggests a tailored sequence of high-level operations to meet the specific needs of the question, leveraging the semantics understanding and reasoning capabilities of LLMs. The core technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method. This method translates the NL question into an Analysis Sketch, which outlines how the table should be prepared to produce the answer, guiding the agent’s reasoning based on this sketch. Compared to the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the semantic relationships between questions and tables.
@@
AutoPrep also includes a set of Programmer agents, each of which synthesizes a question-specific implementation for a given logical data prep operation. However, existing LLM-based code synthesis often generates overly generic code that struggles to effectively address the heterogeneity challenges of tables. For instance, values may have diverse syntactic formats (e.g.,September 1” and “11-24” in $T_{6}$) or semantic representations (e.g., “ITA” and “Italia” in $T_{1}$), making it difficult to generate code tailored to these variations. To address this, we propose a tool-augmented approach that enhances the LLM’s code generation capabilities by incorporating predefined API functions, which allows the LLM to generate more specialized code that accounts for variations in table values. Furthermore, corresponding to the Executing stage, we design an Executor agent that executes the code to process the table.
@@ 2.1. Tabular Question Answering
Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) $\{A_{1},A_{2},\ldots,A_{m}\}$ and $n$ rows $\{r_{1},r_{2},\ldots,r_{n}\}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table. The problem of tabular question answering (TQA) is to generate an answer $Ans$, in response to question $Q$ based on the information in table $T$. By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
@@
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
@@ 2.2. Data Prep for TQA
In contrast to traditional data prep, question-aware data prep for TQA focuses on adapting the table $T$ to the specific informational needs of a given question $Q$, thereby enhancing the semantic alignment between the structured table and the NL question.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$.
As shown in Figure 1, the majority of TQA errors arise from inadequately addressing three key data prep issues: missing semantics, inconsistent values, and irrelevant columns. To address the challenges, we introduce three types of data prep operations.
  • Derive: A data prep task that derives a new column for table $T$ from existing columns, aimed at addressing the challenge of missing semantics. This task typically involves operations such as combining columns through arithmetic computations, extracting relevant values, etc.
  • Normalize: a data prep task that normalizes types or formats of the values in a column of $T$ based on the needs of $Q$, aimed at addressing the challenge of inconsistent values. This task typically involves value representation or format normalization, type conversion, etc.
  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answer question $Q$, aimed at addressing the challenge of irrelevant columns. This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
@@
(a) The Derive operation: Figure 2(a) shows three examples of the Derive operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
(b) The Normalize operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing the formats of Surface, Time and Date for tables $T_{4}$, $T_{5}$ and $T_{6}$ respectively.
(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{7}$ and $T_{8}$. The Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column. Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
Remarks. While issues such as missing values and duplicates are common in real-world datasets, they are significantly less prevalent than the three highlighted challenges in existing TQA benchmarks. Thanks to AutoPrep’s separation of logical operations and physical implementations, the set of data prep operations can be easily extended to accommodate new types of tasks. We leave the exploration of broader data prep challenges to future work.
@@
2.3. A Straightforward LLM-based Solution